Greenhab Report – February 2nd

Crew 221 GreenHab Report 02-02-2020

GreenHab Officers: Marufa Bhuiyan (Crew 220) and Jessica Todd (Crew 221)

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on.

Average Temperature: 21.75°C

Floor Unit: 18°C

Floor Unit: 29°C

Floor Unit: 22°C

Floor Unit: 18°C

Hours of Supplemental Light: 19:30 to 00:00

Daily water usage for crops: 11.1 galleons

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank: 88.44 gallons

Time(s) of watering for crops: 2

07:55= 2.1 galleons

18:40= 9 galleons

Changes to crops: Strawberry plants status unchanged.

Narrative: Today was Sol 0 and the first day in the GreenHab for Crew 221. Today was focused on getting familiar with the GreenHab protocols and procedures, as well as all the beautiful plants we’ll be taking care of for the next two weeks. Crew 221 is ready to demonstrate some serious botany prowess!

Harvest: N/A

Support/supplies needed: N/A

Other Notes:

Strawberry plants are starting to die, please advise.

Aloe and succulents on the upper shelf below the heater are very dry, gave a healthy watering, but will monitor for the next few days, and may need to move them to a cooler location.

Operations Report – February 2nd

Crew 221 Operations Report 2-2-2020

SOL: 0

Name of person filing report: Luis D. Monge

Non-nominal systems: Primary Generator, Hab lower deck 1 fluorescent light ballast (MDRS team in process of upgrade).

Notes on non-nominal systems:


From what time last night: 21:00To what time this morning: 07:00
List any additional daytime hours when the generator was run: 0

Solar— SOC % (Before generator is run at night) 79%

Diesel Reading – 25%

Station Propane Reading – 70%

Ethanol Free Gasoline – gallons

Water (loft tank): 50 gallons
Water Meter: units-
Water (static tank) – (MDRS team full refilling in progress)

Static to Loft Pump used – yes 70 gal

Water in GreenHab –88.44 (MDRS team full refilling in progress)

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging: yes

Spirit rover used: yes
Hours: (before EVA) 128.7
Beginning charge: (Before EVA) 100%
Ending charge: (On return from EVA, before recharging) 85%
Currently charging: yes

Opportunity rover used: no
Hours: n/a
Beginning charge: n/a
Ending charge: n/a
Currently charging: yes

Curiosity rover used:
Hours: 134,8
Beginning charge: 100%
Ending charge: 80%
Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where?

CrewCar used and why, where?

General Notes and Comments: full Cleaning and maintenance performed

Summary of the internet: Nominal.

Summary of Suit and Radios: Suits Nominal

Summary of Hab: Main fluorescent lights in Hab lower deck are being upgraded by the MDRS team. New cameras installed and connected to existing station TV.

Summary of Science Dome operations: Nominal.

Summary of Ram operations: Nominal.

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: MDRS crew all healthy

This is the first ops report for Crew 221. Looking forward to a great time here!

Sol Summary – February 2nd

Crew 221 Sol Summary Report 02-02-2020


Summary Title: Training and preparation time

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary: Crew 221 Earth2Mars arrived on the station meeting members of MDRS-MAU crew 220. The handover of the station was done yesterday from two commanders Shawna and John. Both crews went out for burgers and after a night in the science dome, the crew started their final preparation to start our journey to Mars. Everyone is very excited after one year of preparation to get here to start the rotation.

This morning we waived crew 220 ‘goodbye’, being happy that we could meet this incredible crew and see what they have accomplished. Crew 221 will try its best to also accomplish most of their mission goals.

Today the crew spent all day preparing filling up the water tanks, getting rover training and moving into their new homes and having our first MDRS meals on the completely new oven.

Tomorrow we will wake up on Mars we are looking forward to our launch!

Look Ahead Plan: Internal Crew Training, Voice Protocols, Planning

Anomalies in work: None

Weather: A 20 percent chance of snow before 11 am. Mostly sunny, with a high near -3. Windy, with a northwest wind 45 to 48 km/h.

Crew Physical Status: All crew in good health.

EVA: None

Reports to be filed:

Sol Summary
Photo Report
Operations Report
GreenHab Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site